# Operators with small Kreiss constants

### v1

@@ 5.2. A Pisier-type counterexample
To prove Theorem 1.3, we will adapt a variant of Pisier’s counterexample, given in [Paulsenbook, Chapter 10]. Namely, denote by $S$ the forward shift on $\ell^{2}$: $S(x_{0},x_{1},\dots)=(0,x_{0},x_{1},\dots)$. Next, let $K$ be a Hilbert space and denote by $S_{K}$ the forward shift on $\ell^{2}(K)$: $S_{K}(k_{0},k_{1},\dots)=(0,k_{0},k_{1},\dots)$, whose adjoint acts as $S^{*}_{K}(k_{0},k_{1},\dots)=(k_{1},k_{2},\dots)$. We will take as $F$ the following Foguel-Hankel operator
@@
Here $X=(a_{i+j}W_{i+j})$, where $W_{n}\in\mathcal{B}(K)$ are chosen as in [Paulsenbook, page 142] and $a_{n}$ are complex numbers, whose choice will be specified later. In particular, $\{W_{n}\}$ satisfy the CAR relations modulo the ideal of compact operators. We set $X_{n}=nXS_{K}^{n-1}$ to be the $(1,2)$-entry of $F^{n}$. We summarize the facts which we will use as follows.
Lemma 5.2 ([Paulsenbook]).
@@
Assertion (i) is contained in [Paulsenbook, Theorem 10.5].
The right-hand inequality in (ii) follows from a calculation in the proof of this theorem ([Paulsenbook], page 143). Notice the typo in item (iii) in [Paulsenbook, Theorem 10.5]; it should read as $\sup_{n}n^{2}\sum_{k=n-1}^{\infty}|a_{k}|^{2}<\infty$.
To get the left-hand inequality in (ii), let $\{E_{i,j}\}_{i,j=0}^{\infty}$ be the usual matrix units, considered as linear operators on (scalar) $\ell^{2}$. Set $\widetilde{X}=(a_{i+j}E_{i+j,0})$. Following the proof of Theorem 10.5 in [Paulsenbook], define $\Phi:B(\ell^{2})\to B(K)$ by $\Phi((a_{i,j}))=\sum_{i=0}^{\infty}a_{i,0}W_{i}$. Then $\|\Phi\|\leq 1$. Put $\widetilde{X}_{n}=n\widetilde{X}S^{n-1}$. Then $\Phi(\widetilde{X})=X$ and $\Phi(\widetilde{X}_{n})=X_{n}$, so that we get
@@ 6. Open Questions
Nikolski conjectured [Nikolski] that one should always be able to obtain sublinear growth in $N$ in (1.2) as long as the Kreiss constant remains bounded. In fact, he proved that if $A$ is a diagonalizable $N\times N$ matrix with unimodular eigenvalues, then there exists $\varepsilon>0$ (depending on the basis constant of the eigenvectors of $A$) such that
@@
Unfortunately, it is not clear how exactly $\varepsilon$ depends on $K$ here (see also [Nikolski, Theorem 2.9.9]), while the case of general $N\times N$ matrices currently lies out of reach of Nikolski’s methods. As such, there remains a sizeable gap between the lower bounds (1.5)-(1.6) and Spijker’s estimate (1.2), which naturally motivates the following questions:
@@
Does there exist a continuous function $\varepsilon:(1,\nu_{0})\to(0,\infty)$ such that (1.9) guarantees power-boundedness? What if $\sigma(T)\cap\mathbb{T}=\{1\}?$

### v2

@@ 5.2. A Pisier-type counterexample
To prove Theorem 1.3, we will adapt a variant of Pisier’s counterexample, given in [26, Chapter 10]. Namely, denote by $S$ the forward shift on $\ell^{2}$: $S(x_{0},x_{1},\dots)=(0,x_{0},x_{1},\dots)$. Next, let $K$ be a Hilbert space and denote by $S_{K}$ the forward shift on $\ell^{2}(K)$: $S_{K}(k_{0},k_{1},\dots)=(0,k_{0},k_{1},\dots)$, whose adjoint acts as $S^{*}_{K}(k_{0},k_{1},\dots)=(k_{1},k_{2},\dots)$. We will take as $F$ the following Foguel-Hankel operator
@@
Here $X=(a_{i+j}W_{i+j})$, where $W_{n}\in\mathcal{B}(K)$ are chosen as in [26, page 142] and $a_{n}$ are complex numbers, whose choice will be specified later. In particular, $\{W_{n}\}$ satisfy the CAR relations modulo the ideal of compact operators. We set $X_{n}=nXS_{K}^{n-1}$ to be the $(1,2)$-entry of $F^{n}$. We summarize the facts which we will use as follows.
Lemma 5.2 ([26]).
@@
Assertion (i) is contained in [26, Theorem 10.5].
The right-hand inequality in (ii) follows from a calculation in the proof of this theorem ([26], page 143). Notice the typo in item (iii) in [26, Theorem 10.5]; it should read as $\sup_{n}n^{2}\sum_{k=n-1}^{\infty}|a_{k}|^{2}<\infty$.
To get the left-hand inequality in (ii), let $\{E_{i,j}\}_{i,j=0}^{\infty}$ be the usual matrix units, considered as linear operators on (scalar) $\ell^{2}$. Set $\widetilde{X}=(a_{i+j}E_{i+j,0})$. Following the proof of Theorem 10.5 in [26], define $\Phi:B(\ell^{2})\to B(K)$ by $\Phi((a_{i,j}))=\sum_{i=0}^{\infty}a_{i,0}W_{i}$. Then $\|\Phi\|\leq 1$. Put $\widetilde{X}_{n}=n\widetilde{X}S^{n-1}$. Then $\Phi(\widetilde{X})=X$ and $\Phi(\widetilde{X}_{n})=X_{n}$, so that we get
@@ 6. Open Questions
Nikolski conjectured [22] that one should always be able to obtain sublinear growth in $N$ in (1.2) as long as the Kreiss constant remains bounded. In fact, he proved that if $A$ is a diagonalizable $N\times N$ matrix with unimodular eigenvalues, then there exists $\varepsilon>0$ (depending on the basis constant of the eigenvectors of $A$) such that
@@
Unfortunately, it is not clear how exactly $\varepsilon$ depends on $K$ here (see also [22, Theorem 2.9.9]), while the case of general $N\times N$ matrices currently lies out of reach of Nikolski’s methods. As such, there remains a sizeable gap between the lower bounds (1.5)-(1.6) and Spijker’s estimate (1.2), which naturally motivates the following questions:
@@
Does there exist a continuous function $\varepsilon:(0,\nu]\to(0,\infty)$ such that (1.9) guarantees power-boundedness? What if $\sigma(T)\cap\mathbb{T}=\{1\}?$